OpenStudy (anonymous):

A truck needs 11 gallons of fuel to travel 121 miles. Can the truck travel 99 miles with 9 gallons of fuel? Explain

OpenStudy (chriss):

Yes. The first statement establishes that the truck gets 11 mpg. (121/11=11) Because going 99 miles on 9 gallons of fuel also implies getting 11 mpg, the truck can indeed go 99 miles on 9 gallons of fuel.

OpenStudy (anonymous):

in 1 gallon it travels 121/11 = 11 miles so to travel 99 miles it will need 99/11 = 9 gallons so it can travel 99 miles in 9 gallons of fuel